# Data-driven Error Estimation: Upper Bounding Multiple Errors without Class Complexity as Input

### v3

@@ 2.3 Localizing error bounds
We have introduced error estimation as a data-driven alternative to traditional bounds related to class complexity or size, such as Talagrand’s inequality (Koltchinskii, 2011; Xu & Zeevi, 2020). These methods are frequently enhanced by employing a localization argument to refine the bounds (see, for instance, (Koltchinskii, 2011)). In this section, we apply a similar localization argument to our context to achieve sharper bounds. Suppose that rather than seeking to establish an upper bound on $\max_{h}e_{h}$, we are interested in establishing a bound for the error $e_{h^{*}}$ for the instance $h^{*}:=\operatorname*{arg\,max}_{h}\theta_{h}$  [^8] In fact, $h^{*}$ can be any well-defined instance and possibly depending on $\mathcal{E}$, but independent of $S_{\text{err}}$. For simplicity, we focus on this specific case of $h^{*}:=\operatorname*{arg\,max}_{h}\theta_{h}$ as it illustrates the key idea and is sufficient for our purposes. Extensions to other choices of $h^{*}$ follow similar reasoning. In this case, rather than having to maximize an upper bound across the whole class $\mathcal{H}$, it is enough to maximize across any class that contains $h^{*}$.
To construct such classes, we introduce 2.3 which posits the existence of a lower bound on $\theta_{h^{*}}$. This assumption arises naturally in settings where we have some problem-specific characteristics which provide prior knowledge about the instance with the maximum $\theta_{h}$ value. [^9] For instance, say we are interested in upper bounding the excess risk of a given model in a class (primary goal in Section 3.3). Excess risk is defined as the maximum difference between the risk of the given model and the risk of any other model in the class – hence fits the localization setup. The maximum difference is at least zero and will be positive if the given model is not the best in class – hence satisfies 2.3 with $c=0$. . We will explore concrete examples of this in Section 3.3 and Appendix D.
@@
We have an a priori lower bound on the maximum estimand. Specifically, there exists some constant $c$ such that $c\leq\max_{h}\theta_{h}=\theta_{h^{*}}$.
In this setting, we provide a more general alternative to Theorem 2.2, relying on a weaker alternative to 2.1 (2.4).
@@
Note that while 2.1 relies on the probabilistic bound $\hat{u}(S_{\text{err}},h,\delta)$ to hold for each $h\in\mathcal{H}$, 2.4 only requires a specific probabilistic bound for the instance $h^{*}$. However, since $h^{*}$ is defined in terms of unknown estimands $\theta_{h}$, we cannot directly compute $\hat{u}(S_{\text{err}},h^{*},\delta)$. Instead, our strategy is to use 2.3 to construct a non-increasing sequence of classes $\{\mathcal{H}_{k}(S_{\text{err}})\}_{k\in\mathbb{N}}$ which will all contain $h^{*}$ under certain conditions.
@@
Fix $h^{*}\in\mathcal{H}$ and suppose 2.4 and 2.3 hold for some constant $c$. Further, let $\mathcal{H}_{0}:=\mathcal{H}$ and define $\mathcal{H}_{k}$, $\hat{\xi}_{k}$ inductively as in (7).
$\hat{\xi}_{k}(S_{\text{err}},\delta)$ $:=\max_{h\in\mathcal{H}_{k}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)$ (7)
$\mathcal{H}_{k+1}(S_{\text{err}})$ :=$:=\left\{h\in\mathcal{H}|-\hat{\theta}_{h}\leq\hat{\xi}_{k}(S_{\text{err}},\delta)-c\right\}$
@@
The proof of Corollary 2.5 follows arguments similar to the proof of Theorem 2.2. See Appendix A for details.
@@ 3 Error Estimation Examples
The goal of this section is to better understand error estimation by outlining a simplified recipe for applying error estimation, followed with two examples: (i) bounding error for multiple mean estimates in Section 3.2, and (ii) bounding excess risk in Section 3.3 – which is especially useful when there isn’t a good bound on class complexity.
3.1 Simplifed Recipe for Error Estimation
Let $S_{\text{def}}$ be the dataset used for generating our estimand-estimate-error tuples and $S_{\text{err}}$ be the dataset used for producing upper bounds. In this exposition, we assume for simplicity that the errors we are interested in upper bounding are defined as $e_{h}=\theta_{h}-\hat{\theta}_{h}$, and detail how to construct explicit bounds $\hat{u}(S_{\text{err}},h,\delta)$ to satisfy 2.1.
Notation Explanation
$\theta_{h}$ Estimand for hypothesis $h$
$\hat{\theta}_{h}$ Estimate of $\theta_{h}$ using $S_{\text{def}}$
$e_{h}:=\theta_{h}-\hat{\theta}_{h}$ Error from $S_{\text{def}}$
$\hat{\theta}_{\text{err},h}$ Estimate of $\theta_{h}$ using $S_{\text{err}}$
$e_{\text{err},h}:=\theta_{h}-\hat{\theta}_{\text{err},h}$ Error from $S_{\text{err}}$
Table 1: Key Quantities for Error Estimation Framework. Note that for reasonable estimates, $e_{h}$ and $e_{\text{err},h}$ will be i.i.d. random variables from a symmetric distribution.
@@ 3.1 Simplifed Recipe for Error Estimation
    Generate $S_{\text{err}}$ estimates. In addition to the estimand-estimate-error tuple $(\theta_{h},\hat{\theta}_{h},e_{h})$ from $S_{\text{def}}$, we use $S_{\text{err}}$ to create estimates $\hat{\theta}_{\text{err},h}$ with corresponding errors $e_{\text{err},h}$ (see Table 1). Now, for any $h\in\mathcal{H}$,
@@
    By our construction, $e_{h}$ and $e_{\text{err},h}$ are independent, so an instance $h^{\prime}$ determined by the realization $s$ of $S_{\text{def}}$, can be treated as a single fixed instance in $\mathcal{H}$ with respect to the randomness of $S_{\text{err}}$.
@@
    Construct $\hat{u}(S_{\text{err}},h,\delta)$ to satisfy 2.1 (or 2.4) by applying a single high-probability bound to $e_{\text{err},h}$. In other words, we set
@@
    where $b_{\delta,h}$ is a $1-\delta$ probability upper bound on a single error from $S_{\text{err}}$. Now, conditioning on $s$, under the $(1-\delta)$-probable event $\{e_{\text{err},h^{\prime}}\leq b_{\delta,h^{\prime}}\}$, [^10] Whether we rely on 2.1 or 2.4 depends on whether or not our $b_{\delta,h}$ holds as a probable upper bound for all $h$ or just for $h^{*}$. we have
@@
    Thus we have constructed $\hat{u}(S_{\text{err}},h,\delta)$ which satisfy 2.1, and we can set the max error bound to be $\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$.
@@
    (Optional) Localize the bound. If we are in the setting described in Section 2.3, i.e. bounding $e_{h^{*}}$, where $h^{*}:=\operatorname*{arg\,max}\theta_{h}$, and we have $\theta_{h^{*}}\geq c$ (i.e. 2.3 holds for some constant $c$), we can improve our error bound by Corollary 2.5. We define $\mathcal{H}_{0}=\mathcal{H}$, and $\mathcal{H}_{k}$, $\hat{\xi}_{k}$ as in (7). This gives
@@
    Note that although Corollary 2.5 provides a bound for any $k\in\mathbb{N}$, at some point these upper bounds will converge in practice as $k\to\infty$.
Now under the distribution of $(S_{\text{def}},S_{\text{err}})$, we make a remark tying the distribution of our constructed $\hat{u}(S_{\text{err}},h,\delta)$ (see (10)) to the distribution of $e_{h}$.
Remark 3.1 ($\hat{u}(S_{\text{err}},h,\delta)$ is often a surrogate for $e_{h}$).
Note that in this setting:
@@
Thus, as $e$ and $e_{\text{err}}$ are defined to behave as i.i.d. random variables, our constructed $\hat{u}$ inherits both the distributional and correlational structure of $e$ up to multiplicative and additive constant factors. [^11] If additionally $e$ is symmetrically distributed around the zero vector, then $\hat{u}$ will follow a scaled version of the same distribution. For example, if $e_{h}\sim\mathcal{N}(0,1)$, then $\hat{u}(S_{\text{err}},h,\delta)\sim\mathcal{N}(b_{\delta,h},2)$
Moreover, if $b_{\delta,h}$ is constant across $h\in\mathcal{H}$ (as in examples below),
$\text{Cov}(\hat{u}(S_{\text{err}},h,\delta),\hat{u}(S_{\text{err}},h^{\prime},\delta))=2\text{Cov}(e_{h},e_{h^{\prime}}).$
This indicates that $\hat{u}(S_{\text{err}},h,\delta)$ is a surrogate for $e_{h}$, thus using $\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$ as a bound on $\max_{h\in\mathcal{H}}e_{h}$ is often tight up to constant factors.
@@
In an ideal world, $\hat{u}(S_{\text{err}},h,\delta)$ would be identically distributed as $e_{h}$ (while still being a valid upper bound). Our construction for $\hat{u}(S_{\text{err}},h,\delta)$ often leads to twice as much variance as $e_{h}$ and we need twice as much data to define the same error. Hence, compared to the “ideal” scenario, our error estimation intervals are larger by a factor of $\sqrt{4}=2$.
@@ 3.2 Confidence Interval Construction for Multiple Mean Estimates
Consider a random mapping $Y:\mathcal{H}\rightarrow[-M,M]$ (such as subgroups to average test score), and let $\mathcal{D}$ denote its distribution. Our target estimands are given by $\theta_{h}:=\mathop{\mathbb{E}}_{Y\sim\mathcal{D}}[Y(h)]$ for all $h\in\mathcal{H}$, and our goal is to construct CIs for these estimands that simultaneously hold with high-probability. [^12] Refer to Appendix B for a formal conversion of the subgroup example presented in Section 1 to the setting discussed in this section.We first produce estimates and define errors for each $h\in\mathcal{H}$.
[…]
@@
Table 2: Error-Estimation Quantities for Multiple Means
Thus, for any $h\in\mathcal{H}$, $\hat{\theta}_{h}$ is our mean estimate around which we will construct our CI. The error $e_{h}$ is then defined as the absolute difference between this estimate and the target estimand ($|\theta_{h}-\hat{\theta}_{h}|$) divided by the sample standard error ($\hat{\sigma}_{h}/\sqrt{n}$). Note that upper bounding the error $e_{h}$ gives us an interval for the estimand $\theta_{h}$. Hence, by constructing a high-probability upper bound on the max error ($\max_{h\in\mathcal{H}}e_{h}$), we get simultaneously valid high-probability CIs for all our estimands. The fundamental quantities we establish from dataset $S_{\text{err}}$ are a mean estimate $\hat{\theta}_{\text{err},h}$, sample standard deviation $\hat{\sigma}_{\text{err},h}$, and $e_{\text{err},h}$ (see Table 2). We use these quantities when applying Theorem 2.2 for error estimation.
We now derive an estimate $\hat{u}$ satisfying 2.1. That is, for any $h\in\mathcal{H}$, $\hat{u}(S_{\text{err}},h,\delta)$ should be an upper bound on the input error $e_{h}$ with probability at least $1-\delta$. In this section, we rely on the central limit theorem (CLT) to construct $\hat{u}$ that satisfies this guarantee asymptotically (a sample size of $n\geq 30$ is often considered sufficient for the asymptotic guarantees of CLT to hold). [^13] Our choice to rely on CLT here is due to its popularity in practice. To get non-asymptotic results, one may construct $\hat{u}$ based on Hoeffding or Bernstein’s inequality. By construction, we have defined $e_{h}$ and $e_{\text{err},h}$ to be distributed as $|\mathcal{N}(0,1)|$ random variables. Thus, from the CLT, $z_{\delta/2}$ upper bounds $e_{\text{err},h}$ asymptotically with probability at least $1-\delta$, where $z_{\delta}$ is the $\delta$ critical value for the standard normal distribution ($N(0,1)$). [^14] That is, a sample from the standard normal distribution is greater than $z_{\delta}$ with probability $\delta$. Now,
$e_{\text{err},h}\leq z_{\delta/2}$ (12)
$\implies$ $e_{h}=\frac{\sqrt{n}|\hat{\theta}_{h}-\theta_{h}|}{\hat{\sigma}_{h}}$
@@
Hence, for all $h$, $\Pr_{S_{\text{err}}}(\hat{u}(S_{\text{err}},h,\delta)\geq e_{h}|S_{\text{def}})\geq 1-\delta$ asymptotically holds. Thus when $n$ is large enough for CLT-based CIs to hold, we have 2.1 holds. Hence from Theorem 2.2, we have the following corollary.
@@
Consider the setup in Section 3.2. Suppose the following CLT-based CIs hold: $\Pr_{S_{\text{err}}}(\hat{u}(S_{\text{err}},h,\delta)\geq e_{h}|S_{\text{def}})\geq 1-\delta$ for all $h\in\mathcal{H}$. Then for $\hat{\xi}(S_{\text{err}},\delta):=\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$, we have (13) holds.
@@
Proof.
From Theorem 2.2, we have $\Pr_{\bar{S}}\left(\max_{h\in\mathcal{H}}e_{h}\leq\hat{\xi}(S_{\text{err}},\delta)\right)\geq 1-\delta$. The result follows from noting that $\max_{h\in\mathcal{H}}e_{h}\leq\hat{\xi}(S_{\text{err}},\delta)$ is equivalent to $\forall h\in\mathcal{H},\;|\hat{\theta}_{h}-\theta_{h}|\leq\frac{\hat{\sigma}_{h}}{\sqrt{n}}\hat{\xi}(S_{\text{err}},\delta)$. ∎
@@
Remark 3.4 (Tightness compared to union bound).
Our error estimation bound is the maximum of $\hat{u}(S_{\text{err}},h,\delta)$, to further bound this analytically we can use the union bound. Under the randomness of $\bar{S}$, one can show that for the multiple mean setting $\hat{u}(S_{\text{err}},h,\delta)$ is asymptotically distributed as $z_{\delta/2}+|N(0,2)|$. Applying the union bound, we get our error estimation bound is further bounded by $O(\sqrt{\log|\mathcal{H}|})$ – showing that, up to constant factors, error estimation is no worse than applying a direct union bound on errors. [^15] For the multiple mean setting, $e_{h}$ is asymptotically distributed as $|N(0,1)|$. Further, these bounds are tight (up to constants) when our mean estimation tasks in $\mathcal{H}$ are independent. Thus, although our approach doesn’t rely on $|\mathcal{H}|$ as an input, our bound may still grow as a function of $|\mathcal{H}|$. In practice, our estimation tasks are not independent, and in such cases our method can provide significantly tighter bounds by escaping the explicit dependence on complexity or size of $\mathcal{H}$.
Before moving on to the infinite class setting, we visualize an example of this tightness in Figure 1 where we errors are correlated, and we see that as $|\mathcal{H}|$ grows, the union bound becomes progressively looser, while our error bound remains stable. We also see that for smaller group sizes $|\mathcal{H}|$ our bound may be looser, as we pay a constant factor price (see Remark 3.2).
Figure 1: In this plot, we bound maximum error across estimates quantile estimates. We use $n=10000$ i.i.d. samples from a standard normal distribution to estimate an increasing number of quantiles. We use two methods to bound the maximum error, and plot the excess (bound - true maximum error) of error bounds vs. increasing number of quantiles estimated for each: traditional union bound (orange) and our error estimation method (blue).

### v4

@@ 2.3 Localizing error bounds
Traditional complexity bounds are frequently enhanced by employing a localization argument by bounding over a subclass near the optimum (Koltchinskii, 2011). We apply a similar argument to data-driven bounds. Suppose that rather than bounding $\max_{h}e_{h}$, we are interested in an upper bound for the error $e_{h^{*}}$ for the instance $h^{*}:=\operatorname*{arg\,max}_{h}\theta_{h}$  [^7] In fact $h^{*}$ can be any fixed instance (possibly depending on $\mathcal{E}$) independent of $S_{\text{err}}$. We focus on $h^{*}:=\operatorname*{arg\,max}_{h}\theta_{h}$ for clarity, other choices follow similar reasoning. In this case, rather than having to maximize an upper bound across the whole class $\mathcal{H}$, it is enough to maximize across any class that contains $h^{*}$.
To construct such classes, we introduce 2.3 which posits the existence of a lower bound on $\theta_{h^{*}}$. This assumption arises naturally in settings where problem-specific characteristics provide prior knowledge. [^8] For instance, in Section 3.3, excess risk is the maximum difference between the risk of the given model and the risk of any other model in the class – hence fits the localization setup of 2.3 with $c=0$.
@@
We have an a priori lower bound on the maximum estimand. Specifically, there exists some constant $c$ such that $c\leq\max_{h}\theta_{h}$.
In this setting, we base our result on a weaker alternative to 2.1 (2.4).
@@
While 2.1 relies on the probabilistic bound $\hat{u}(S_{\text{err}},h,\delta)$ to hold for each $h\in\mathcal{H}$, 2.4 only requires a specific probabilistic bound for the instance $h^{*}$. However, since $h^{*}$ is defined in terms of unknown estimands $\theta_{h}$, we cannot directly compute $\hat{u}(S_{\text{err}},h^{*},\delta)$. Instead, our strategy is to use 2.3 to construct a non-increasing sequence of classes $\{\mathcal{H}_{k}(S_{\text{err}})\}_{k\in\mathbb{N}}$ which will all contain $h^{*}$ under certain conditions. We arrive at the following corollary, see Appendix D for detailed proof.
@@
Fix $h^{*}\in\mathcal{H}$ and suppose Assumptions 2.3 and 2.4 hold for some constant $c$. Further, let $\mathcal{H}_{0}:=\mathcal{H}$ and define $\mathcal{H}_{k}$, $\hat{\xi}_{k}$ inductively as in (5).
$\hat{\xi}_{k}(S_{\text{err}},\delta)$ $=\max_{h\in\mathcal{H}_{k}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)$ (5)
$\mathcal{H}_{k+1}(S_{\text{err}})$ $=\left\{h\in\mathcal{H}|-\hat{\theta}_{h}\leq\hat{\xi}_{k}(S_{\text{err}},\delta)-c\right\}$
@@ 3 Error Estimation Examples
The goal of this section is to better understand error estimation. We outline a simplified recipe (Section 3.1), followed by two examples: bounding error for multiple mean estimates (Section 3.2), and bounding excess risk in supervised learning tasks (Section 3.3). The latter showcases how error estimation can replace complexity-based bounds, and how localization can further refine them.
3.1 Simplified Recipe for Error Estimation
Let $S_{\text{def}}$ and $S_{\text{err}}$ be the error defining and error estimating datasets respectively. For simplicity of exposition, we focus on errors of the form $e_{h}=\theta_{h}-\hat{\theta}_{h}$, and detail how to construct explicit bounds $\hat{u}(S_{\text{err}},h,\delta)$ to satisfy 2.1.
Table 1: Key quantities used for an estimand $\theta_{h}$.
Defining data $S_{\text{est}}$ Error estimation data $S_{\text{err}}$
Estimate $\hat{\theta}_{h}$ $\hat{\theta}_{\text{err},h}$
Error $e_{h}=\theta_{h}-\hat{\theta}_{h}$ $e_{\text{err},h}=\theta_{h}-\hat{\theta}_{\text{err},h}$
@@ 3.1 Simplified Recipe for Error Estimation
    Generate $S_{\text{err}}$ estimates. In addition to the tuple $(\theta_{h},\hat{\theta}_{h},e_{h})$ from $S_{\text{def}}$, we use $S_{\text{err}}$ to define estimates $\hat{\theta}_{\text{err},h}$ and errors $e_{\text{err},h}$ (see Table 1). Now, for each $h\in\mathcal{H}$,
@@
    By construction, $e_{h}$ and $e_{\text{err},h}$ are independent, so an instance $h^{\prime}$ determined by a realization of $S_{\text{def}}$, can be treated as a single fixed instance in $\mathcal{H}$ with respect to the randomness of $S_{\text{err}}$.
@@
    Construct $\hat{u}$ to satisfy 2.1 (or 2.4) by applying a single high-probability bound to $e_{\text{err},h}$. If $b_{\delta,h}$ is a $1-\delta$ probability upper bound on $e_{\text{err},h}$, we set
@@
    Now, conditioning on $s$, under the $(1-\delta)$-probable event $\{e_{\text{err},h^{\prime}}\leq b_{\delta,h^{\prime}}\}$, we have
@@
    Thus we satisfy 2.1, and can set the max error bound $\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$.
@@
    (Optional) Localize the bound. In the setting from Section 2.3, we can improve our error bound by Corollary 2.5. We define $\mathcal{H}_{0}=\mathcal{H}$, and $\mathcal{H}_{k}$, $\hat{\xi}_{k}$ as in (5). This gives
@@
    In practice, the bounds $\hat{\xi}_{k}$ converge as $k\to\infty$.
We make a couple of remarks about our resulting method.
Remark 3.1 (Relationship between $\hat{u}$ and e).
By construction:
@@
Thus, when $e$ and $e_{\text{err}}$ are i.i.d. random variables, $\hat{u}$ inherits a closely related distribution to ee. [^9] For example, if $e_{h}\sim\mathcal{N}(0,1)$, then $\hat{u}(S_{\text{err}},h,\delta)\sim\mathcal{N}(b_{\delta,h},2)$ Moreover, $\hat{u}$ preserves the dependence pattern across instances $h\in\mathcal{H}$. In particular, if $b_{\delta,h}$ is constant across $h\in\mathcal{H}$ (as in examples below), then $\text{Cov}(\hat{u}(S_{\text{err}},h,\delta),\hat{u}(S_{\text{err}},h^{\prime},\delta))=2\text{Cov}(e_{h},e_{h^{\prime}})$. Thus $\hat{u}(S_{\text{err}},h,\delta)$ is a surrogate for $e_{h}$, so $\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$ often closely tracks $\max_{h\in\mathcal{H}}e_{h}$.
@@
In an ideal world, $\hat{u}(S_{\text{err}},h,\delta)$ would be identically distributed as $e_{h}$ (while still being a valid upper bound). Our construction for $\hat{u}(S_{\text{err}},h,\delta)$ often leads to twice as much variance as $e_{h}$ and we need twice as much data to define the same error. Hence, compared to the “ideal” scenario, our error estimation intervals are larger by a factor of $\sqrt{4}=2$. Cross-fitting can substantially reduce this sample-splitting inflation while retaining validity; see Appendix G.
Remark 3.3 (Tightness compared to union bound).
Since $\hat{u}(S_{\text{err}},h,\delta)$ is a set of random quantities, a union bound over $h\in\mathcal{H}$ can be applied to (loosely) bound $\max_{h}\hat{u}(S_{\text{err}},h,\delta)$. Thus, in finite cases, our method is always no worse than the union bound up to constant factors. In practice, when errors are correlated, our method can provide significantly tighter bounds (see Figure 1).
Figure 1: Error estimation vs. union bound. As correlation between tasks increases, the true max error (green) decreases. Our bound (orange) adapts accordingly while the union bound (blue) stays fixed due to its explicit dependence on $|\mathcal{H}|$. See Appendix C for details.
@@ 3.2 Confidence Interval Construction for Multiple Mean Estimates
Let $Y:\mathcal{H}\rightarrow[-M,M]$ be a bounded random map (such as subgroups to average test score), and define $\theta_{h}:=\mathop{\mathbb{E}}_{Y\sim\mathcal{D}}[Y(h)]$ for all $h\in\mathcal{H}$. Our goal is to construct simultaneously valid CIs for these estimands that hold jointly with probability $1-\delta$. [^10] See Appendix E for a conversion of Example 1 to the setting in Section 3.2. We draw $2n$ i.i.d. samples, and split them randomly into two independent datasets of equal size (for simplicity), $S_{\text{def}}$ and $S_{\text{err}}$. We define central quantities in Table 2.
Table 2: Error-Estimation Quantities for Multiple Means
@@
For any $h\in\mathcal{H}$, $\hat{\theta}_{h}$ is our mean estimate around which we will construct our CI. The error $e_{h}$ is the absolute difference between this estimate and the target estimand ($|\theta_{h}-\hat{\theta}_{h}|$) divided by the sample standard error ($\hat{\sigma}_{h}/\sqrt{n}$). By constructing a high-probability upper bound on the max error ($\max_{h\in\mathcal{H}}e_{h}$), we get simultaneously valid high-probability CIs for all our estimands. We establish $\hat{\theta}_{\text{err},h}$, $\hat{\sigma}_{\text{err},h}$, and $e_{\text{err},h}$ from dataset $S_{\text{err}}$ (see Table 2).
In this section, to derive an estimate $\hat{u}$ satisfying 2.1, we implement a central limit theorem (CLT)-based asymptotic guarantee. [^11] We use CLT due to its popularity, however, non-asymptotic alternatives, such as Hoeffding or Bernstein’s inequality would lead to finite-sample guarantees. By construction, we have defined $e_{h}$ and $e_{\text{err},h}$ to be distributed as $|\mathcal{N}(0,1)|$ random variables. Thus, from the CLT, $z_{\delta/2}$ upper bounds $e_{\text{err},h}$ asymptotically with probability at least $1-\delta$, where $z_{\delta}$ is the $\delta$ critical value for the standard normal distribution ($N(0,1)$). Now, if indeed $e_{\text{err},h}\leq z_{\delta/2}$ holds, then
$e_{h}$ $=\frac{\sqrt{n}|\hat{\theta}_{h}-\theta_{h}|}{\hat{\sigma}_{h}}$ (10)
@@
Hence, for all $h$,
$\Pr_{S_{\text{err}}}(\hat{u}(S_{\text{err}},h,\delta)\geq e_{h}|S_{\text{def}})\geq 1-\delta$
asymptotically holds, so for $n$ is large enough for CLT-based CIs to hold, we have satisfied 2.1. From Theorem 2.2, we have the following corollary.
@@
Consider the setup in Section 3.2. Suppose the following CLT-based CIs hold:
$\Pr_{S_{\text{err}}}(\hat{u}(S_{\text{err}},h,\delta)\geq e_{h}|S_{\text{def}})\geq 1-\delta$
for all $h\in\mathcal{H}$. Then for $\hat{\xi}(S_{\text{err}},\delta):=\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$, we have (11) holds.